OpenStudy (anonymous):

a circle has a radius of 12. find the diameter and area. a. r=24 a= 9pie b. r= 24 a=36pie c. r= 6 a=36pie

OpenStudy (agent47):

diameter = 2r radius=pi*r^2

OpenStudy (agent47):

it's none of the options

OpenStudy (agent47):

diameter is 24 area is 144pi

OpenStudy (anonymous):

Sorry, the r's should be d's.

OpenStudy (anonymous):

radius of 12 so, diameter= 2*radius=2*12=24

OpenStudy (anonymous):

area= pi*radius^2= pi*12^2= 144pi